### v1

@@ 5 Discussion
(a) Glyph dataset.
(b) OSM dataset.
Figure 8: Failure cases of PolyMP on Glyph and OSM datasets.
@@
Figure 9: Feature maps to visualize salient features learned by individual models to classify input geometries to output categories. The darker the color, the more salient are the features learned. Feature values are normalised $\in[0,1]$.
@@ 5.1 Model Expressivity
For the interpretation of model expressivity with respect to various polygon encodings and learning models, we visualize the feature maps of models tested on original and geometric transformed samples in glyph dataset in Figure 9.
Figs. 9a and 9b depict the salient feature maps and class predictions of four models (Deepset, Set Transformer, GCAE and PolyMP) on geometric transformed samples (sheared and rotated glyph M). In this case, the PolyMP correctly predicts the class label of glyph $M$, whereas DeepSet, Set Transformer and GCAE have made incorrect predictions on the geometric transformed samples. By comparing the feature maps in Figures 9c and 9d, the PolyMP learns to identify the global structures or the “skeleton” of polygons via non-trivial vertices on the boundary. The message-passing network and local pooling function (i.e., max-pooling) can capture the salient vertices from local neighbours (i.e., non-trivial vertices (dark color) in the clusters of trivial vertices (light color), which enable a robust embedding of polygons under geometric transformations. Set-based models and GCAE learn sparser geometric features for input polygons, capturing salient features regionally in node clusters and ignoring geometric features of more distant nodes. This can be attributed to the set representation where any relational information between nodes is neglected.
Figures 9e and 9f demonstrate the feature maps on polygons with holes. Both Deepset and Set Transformer only capture geometric features from non-trivial vertices located at the exterior boundary of polygons. While set-based models accept point set encoding of polygons in varying length and the output features are invariant to the feed-in order of vertices, such methods largely ignore the connectives and interactions between points. Such a learning setting inherently hinders the model’s capability to learn geometric information from polygons with holes. The graph convolution model, GCAE, lacks learning flexibility in learning local geometric features as discussed in the previous sections. On the other hand, PolyMP can capture geometric features from salient points located at the exterior and interiors of polygons. In the message-passing network, the exterior and interior of polygons are considered to be two distinct elements as there is no direct connection between points located at the two boundaries. In other words, the geometric features (i.e., the “message”) are only passed between points that are connected in a connected component of a graph, in this case defined by the linear rings of polygon boundaries.
@@ 5.2 Limitations
Of course, PolyMP still encounters limitations in terms of misclassification. Fig. 8. demonstrates failures of PolyMP classifications on the Glyph and OSM datasets. On synthetic data, misclassifications occur predominantly on samples where excessive geometric transformations (e.g., heavy shearing and anisotropic scaling) destroy the gestalt of the shape to the extent it is no longer label preserving (Fig. 8(a): letter V, Z and R). A special case are shapes with different labels dependent on orientation – M and W, where depending on the extent of rotation and shearing, misclassifications occur. For instance in row two of Fig. 8(a), we observe that rotated letters W can be misclassified into letter M, or even Z; heavily sheared W can be misclassified as H and V. Similarly for building polygons (OSM), PolyMP may misclassify polygons sharing similar geometric shapes (E, vs L or U) (Fig. 8).
Considering the characteristics of the two datasets (Sec. 3.1), PolyMP trained Glyph dataset learns locally salient geometric features via message-passing layers from serif glyphs, which contain salient decorative strokes ending. Letters sharing similar geometric shapes are misclassified when the geometric shapes of classified polygons are destroyed by excessive geometric transformations.
@@ 6 Conclusion
Based on extensive empirical evidence given by the analysis of invariances to geometric transformations, trivial vertices and model generalizability, we demonstrate that discrete, non-grid like data representations in combination with graph-based message-passing neural networks (PolyMP) are optimal for deep learning models on 2D vector geometries. We demonstrate a robust ability to learn shapes of polygonal geometries, and demonstrate the ability to generalize to a real-world building footprint dataset.
We find that a learning model based on graph data representation that explicitly encodes both node and explicit edge structures leads to a robust approach enabling highly accurate feature learning on geometric shapes – conveniently, these structures are readily available from the 2D simple feature representations of geometries implemented in current GIS libraries. Our results indicate that representations play an important part providing a strong inference bias for geometric shape learning, and tasks such as the classification of polygons and boosting model performance. We believe that our approach to learning geometric-invariant features of polygons is not limited to simple geometric shape classification. This approach can benefit suitable geospatial tasks where the geometric-invariant features of polygons matches the learning objectives (i.e., geospatial regression tasks involving spatial relation predictions of polygon map [8]).
We also discuss the limitations of PolyMP through misclassification examples on datasets. Through the failure cases of classification, we observe that the message-passing network for graph data representation of polygons has limited capability of transferring locally salient geometric features from one domain to another. The simple message-passing network has limited control on learning/not learning geometric invariant features as need.
The geometric invariant feature learning of vector geometries is currently an area of intense investigation in spatial deep learning [2, 1, 6, 35, 8]. Unlike in learning based on raster datasets or point clouds, the segmentation problem is partially solved by the explicit encoding of structured vector geometries. Here, we show results that allow us to undertake the classification of spatial geometric shapes efficiently, based on decisions that can be grounded in theory and are experimentally validated. In future work, we will further examine how richer graph encoding of the structure of geometries, as well as of the relationships between proximal geometries could further boost model performance, or enable the segmentation of complex spatial object arrangements.
Statements and Declarations
Data availability. The data and codes that support the findings of this study are available with the identifier(s) at the private link: https://figshare.com/s/b781d7724d3220e24c7c.
Conflicts of interest. This paper has been approved by all co-authors. The authors have no competing interests to declare that are relevant to the content of this article.

### v2

@@ 6 Discussion
Figure 8: Feature maps to visualize geometric features learned by individual models to classify input geometries to output categories. The darker the color, the more salient are the features learned. Feature values are normalised $\in[0,1]$.
(a) Glyph dataset.
(b) OSM dataset.
Figure 9: Failure cases of PolyMP on Glyph and OSM datasets.
@@ 6.1 Model Expressivity
To interpret the expressivity of different polygon encoding and learning models, we visualize the feature maps of models tested on original and geometrically transformed samples from the Glyph dataset in Figure 8.
Figures 8a and 8b illustrate the salient feature maps and class predictions of set-based (DeepSet, Set Transformer) and graph-based (GCAE, DSC-NMP, PolyMP, and PolyMP-DSC) learning models on geometrically transformed samples. Notably, PolyMP and PolyMP-DSC correctly predict the class label of the sheared and rotated glyph $M$, whereas DeepSet, Set Transformer, GCAE, and DSC-NMP make incorrect predictions. By comparing the feature maps in Figures 8c and 8d, we observe that PolyMP and PolyMP-DSC effectively capture the global structure or “skeleton” of polygons by identifying salient, non-trivial vertices along the boundary. The message-passing encoder, coupled with a local pooling function (i.e., max-pooling), extracts geometric features from local neighborhoods, where non-trivial vertices (darker color) are distinguished within clusters of trivial vertices (lighter color). This enables the models to generate robust embeddings of polygons, even under geometric transformations. In contrast, set-based models and GCAE learn sparser geometric representations, capturing regional features in node clusters while neglecting relational information between more distant nodes. This limitation arises from their reliance on set representations, which inherently discard connectivity information between vertices.
Figures 8e and 8f illustrate the feature maps for polygons with holes. DeepSet and Set Transformer primarily capture geometric features from non-trivial vertices along the exterior boundary, while largely ignoring those within the interior. Although set-based models can encode polygons of varying lengths and produce permutation-invariant feature representations, their lack of connectivity modeling restricts their ability to learn geometric structures in polygons with holes. Similarly, the graph convolution model GCAE struggles to flexibly capture local geometric features, as discussed in previous sections. In contrast, PolyMP and PolyMP-DSC effectively capture geometric features from both the exterior and interior boundaries of polygons. This is achieved through graph message-passing encoders, where geometric features (i.e., the “message”) are exchanged only between directly connected nodes within the same connected component of a graph—defined in this case by the polygon’s linear rings. As a result, PolyMP-based models can better preserve structural relationships in polygons with complex topologies.
@@ 6.2 Limitations
Despite its strong performance, our proposed method still encounters misclassification limitations. Figure 9 highlights instances where PolyMP fails to classify samples correctly in the Glyph and OSM datasets.
On synthetic data, misclassifications predominantly occur in cases where excessive geometric transformations—such as heavy shearing and anisotropic scaling—severely distort the shape’s structure, rendering it unrecognizable: rotated W shapes (Fig.9a) are sometimes mistaken for M or even Z, while heavily sheared W shapes may be misclassified as H or V.
Similarly, in the OSM dataset, PolyMP and PolyMP-DSC occasionally misclassify building polygons with similar geometric structures, such as E, L, and U shapes (Fig. 9b). This can be attributed to the model’s reliance on local geometric features, which may not sufficiently distinguish between structures with high shape similarity.
Considering the characteristics of the two datasets (Sec. 4.1 and 4.2), PolyMP and PolyMP-DSC trained on the Glyph dataset primarily learns locally salient geometric features through message-passing encoders. These features are particularly effective for serif glyphs, which contain distinctive decorative stroke endings. However, when excessive geometric transformations distort these key features, the model struggles to preserve label consistency, leading to misclassification among shapes with similar structures.
@@ 7 Conclusion
In this study, we investigated the challenge of geometric-invariant shape classification from 2D vector polygons, a common but underexplored data format in spatial analysis. Through comprehensive empirical evaluation, we demonstrated that combining discrete, non-grid-like polygon representations with graph-based message-passing neural networks—PolyMP and its densely self-connected variant, PolyMP-DSC—enables the learning of robust and expressive geometric features.
Our findings show that graph representations of polygons, which encode both geometric structure and vertex connectivity, significantly outperform sequence- and set-based encodings in terms of transformation invariance and generalization. In particular, PolyMP-DSC enhances feature propagation by incorporating dense self-connections, resulting in improved classification performance and robustness to structural perturbations such as trivial vertices. These properties are crucial for geospatial tasks where object shapes may undergo rotation, scaling, shearing, or digitization noise.
Importantly, our approach aligns well with practical geospatial workflows: polygon-to-graph conversion is natively supported by many GIS libraries, enabling seamless adoption of our models in downstream applications such as automated cartographic generalization, building footprint recognition, and road geometry analysis.
Nevertheless, current limitations remain. Both PolyMP and PolyMP-DSC occasionally struggle with locally distinctive features that vary across shape classes or domains. To address this, future work will explore the integration of complementary geometric-invariant descriptors—such as spectral features (e.g., NUFT) and local shape attributes (e.g., turning angles, curvature, or radii of arc segments)—to further improve model robustness and invariance.
Overall, this work highlights the importance of representation choice in spatial deep learning. By designing models with inductive biases that reflect geometric structure and transformation invariance, we move closer to human-like perception of shape in geospatial domains. Beyond classification, our findings open promising directions for extending graph-based learning to polygon regression, spatial clustering, and topological inference in large-scale vector datasets.
Declarations
Data availability.
The data and code supporting the findings of this study are available at https://github.com/zexhuang/PolyMP.
Conflicts of interest.
This paper has been approved by all co-authors. The authors have no competing interests to declare that are relevant to the content of this article.
Funding.
No funding was obtained for this study.